OpenStudy (anonymous):

Divide Rational Expression and put in Simplest form? 24y^2 ÷ 18y x^2+16x+64 x^2 +8x

OpenStudy (campbell_st):

you 1st question can be written \[\frac{24y^2}{18y} = \frac{4\times6\times y \times y}{3\times6\times y}\] cross out the common factors... to get you answer

OpenStudy (anonymous):